6. An isosceles triang has perimet,er 30 cm and each of the equal sides is 12 cm. Find the area of the triangle.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Each equal side of the isosceles triangle is 12 cm. :. Third side = (30 - 12 - 12) cm = 6 cm. So, sides of triangle are 12 cm, 12 cm and 6 cm. 12 + 12 + 6 s 2 = 15 Now, area = 15(15 - 12X15 - 12)(16 - 6) = 16 x 3 x 3 x 9 = 9 .J1r, cm2